OpenStudy (anonymous):

"A 14-foot ladder is leaning against a house with the base of the ladder 5 feet from the house. How high up the house does the ladder reach? If necessary, round to the nearest tenth foot." Is it 9ft???

OpenStudy (amistre64):

how level is the ground? and how level is the house?

OpenStudy (amistre64):

h = sqrt(14^2 - 5^2) in a perfect world :)

OpenStudy (amistre64):

169 25 ---- 144 might be 12

OpenStudy (anonymous):

h = V ( 196 - 25 ) = 13.1 ft

OpenStudy (amistre64):

196 25 ---- 171 bigger than 12